# Comparative Insights on Adversarial Machine Learning from Industry and Academia: A User-Study Approach

### v1

@@ Abstract
An exponential growth of Machine Learning and its Generative AI applications brings with it significant security challenges, often referred to as Adversarial Machine Learning (AML). In this paper, we conducted two comprehensive studies to explore the perspectives of industry professionals and students on different AML vulnerabilities and their educational strategies. In our first study, we conducted an online survey with professionals revealing a notable correlation between cybersecurity education and concern for AML threats. For our second study, we developed two CTF challenges that implement Natural Language Processing and Generative AI concepts and demonstrate a poisoning attack on the training data set. The effectiveness of these challenges was evaluated by surveying undergraduate and graduate students at Carnegie Mellon University, finding that a CTF-based approach effectively engages interest in AML threats. Based on the responses of the participants in our research, we provide detailed recommendations emphasizing the critical need for integrated security education within the ML curriculum.
@@ I Introduction
The modern-day computing ecosystem consists of a large volume and variety of data. This data can be viewed in different ways by different professionals. For cybersecurity professionals, data is primarily seen through the lens of security, encompassing access levels, authorization, and data integrity. In contrast, machine learning professionals approach data as a resource, primarily for analysis, focusing on datasets and models. To explore the similarities and differences in these views, as well as to address various aspects of security, privacy, and machine learning models in today’s world, we conducted comprehensive user studies. By capturing and analyzing these different perceptions, our goal was to bridge the gap between industry practices and academic research for AML threats, while developing a CTF based educational approach for Adversarial Machine Learning (AML).
@@
In this paper, we undertake a thorough exploration of AML, examining various aspects which include theory on AML attacks and their implications, expanding our knowledge of AML attack to practical prompt-based learning models, a Capture-the-Flag (CTF) perspective on AML, versus an industry perspective of AML. In this research, we explore different aspects of issues in Generative AI through the lens of two user studies and derive a novel educational approach for students and professionals for a defense strategy against AML attacks.
The first study involves an online survey of $12$ cybersecurity and machine learning experts where we collected and analyzed their experiences and perspectives on AML trends and defenses in the industry. In our second study, we used the concept of Generative AI to construct two Capture-the-Flag (CTF) challenges which allowed us to evaluate two perspectives—develop our proof of concept for ML security CTF challenges for the picoCTF platform (a widely used university-hosted Capture-the-Flag educational website) and evaluate CTF as a mode of education for AML threats. In the same study, we also developed a CTF challenge where the user has to utilize Feature Collision and Convex Polytope attacks against a Generative AI chat bot with data loopback to influence the training data.
@@
  • We developed and evaluated our novel Capture-the-Flag (CTF) challenges that are a proof-of-concept implementation of AML concepts, thus providing a practical and engaging learning approach for students.
@@ II-A Taxonomy of AML Attacks
Huang et al. [22] categorizes AML attacks based on three properties:
@@
  • Oracle Attacks: In this type of attack, adversaries exploit the input-output pairings to gather information on the model like the parameters (Extraction attack), reconstruct training data (Inversion attacks) and determine in a brute force manner whether any of the supplied points were part of training data set (Membership Interference Attack).
@@ II-B User-Centric Perspective of AML
Although there are various AML defense strategies and related research, we believe it is imperative to explore the impact and cause of human perception on the efficacy of these strategies. Not only does this point of view help devising effective defenses, it also helps us understand the best technique of learning for the people who are actually involved in this line of work.
@@
While the aforementioned studies give a deeper insight into AML and the need for its education, there seems to be little light on comparison of perspectives of two pivotal entities of the field: the professionals and students in the fields of Cybersecurity and Machine Learning. Hence, our aim is to identify the current attack landscape in the industry and associated challenges faced by students and professionals in detecting and mitigating AML attacks. We also aim to explore an academic perspective by developing AML CTF challenges in order to understand the student’s point of view for this new concept in cybersecurity.
@@ II-C Defenses against AML
In this section, we review prior research on defensive strategies against Adversarial Machine Learning (AML) threats to understand the current security framework for ML systems. Goodfellow et al. [11] theorize a “Generative Adversarial Networks (GAN)” framework for estimating generative models using adversarial networks, and the authors proposed to train two models:
@@ III-A1 Ethical Considerations
Our data collection process for our study was organized in such a way that ethical considerations were taken into account in each step of the way. Our survey script, acknowledgments, recruitment materials, and other relevant documents were approved by our institution’s Institutional Review Board (IRB). The participants were compensated 5 USD for the 10 minute online survey, which is equivalent to 30 USD per hour, which is well above the minimum wage.
@@ III-A2 Recruitment
We conducted an online survey of 12 participants who are working professionals with varied backgrounds in Security, Machine Learning, and Privacy. To streamline participant selection, we utilized two platforms: WiCyS (Women in Cybersecurity Conference) 2023 [38] where participants were recruited through posters distributed to attendees, and an online platform connecting students of our university with their alumni where we post our recruitment text on a message board. The interested participants in both cases were asked to contact us by email. After receiving 24 inquiries collectively from both platforms, we evaluated participants’ backgrounds to verify that they were above the age of 18 and that they are a working professional in either ML or security, and selected 12 qualifying participants out of them. The survey was hosted on Qualtrics [31] and each participant was sent a unique link to record their responses.
@@ III-B Analysis
Our survey was taken by 12 participants. About 64% of the participants belonged to an age ranging between 25 and 34 years. The data collected was checked for possible errors, such as missing or irrelevant text entries. We also checked responses to see if any participants have zero experience in all three fields—Cybersecurity, Privacy, and Machine Learning. After finding no errors, we analyzed the study data as listed below.
@@ III-B1 Background of Participants
Our understanding of the correlation between the participants’ background in Machine Learning, Cybersecurity and Privacy and their concern for security in Artificial Intelligence and Machine Learning led us to ask them non-personally identifiable information such as participant’s age, gender, education level and relevant experience in the above mentioned fields. Regarding the background of the participants, we analyzed two key metrics, namely “Education” and “Work Experience”. The table below shows the level of education and the level of experience that correspond to the number of participants in the given category. This table (Table I) shows a categorization as relevant to our hypothesis by aggregating some results. The more detailed breakdown is available in the Appendix.
@@
As seen in the table, the participants in this study display a diverse background and a set of qualified skills, demonstrating proficiency in matters related to security and privacy issues in Machine Learning models. About 95% of the participants have worked with the ML algorithms in some form—Regression models (9 out of 12), Decision tree (10 out of 12), Support Vector Machines (6 out of 12), Neural Networks (9 out of 12) and Other with includes Clustering algorithms (1 out of 12) and reinforcement learning (1 out of 12). Based on these observations, we explore our first hypothesis:
@@
From the above table, we observe an “expertise effect” for participants having a background in educational or work experience in cybersecurity since their backgrounds are correlated with their concern for AML threats, and hence are more likely to think about security in ML. However, a background in ML itself does not show any correlation with the participant’s concern for security issues in ML. Lastly, an educational background in privacy is correlated with the participant’s concern, whereas merely work experience is not. This result highlights the importance of security and privacy education amongst professionals, especially ML engineers.
@@ III-B2 Personal Experience in the industry
In the studies of Hickman et al. [20] and Rees et al. [32], the authors reviewed official documents, which consisted of EU guidelines for Trustworthy AI and organizational and government documents, respectively. In both cases, the authors have concluded that while the documents aim to address the concerns in AI, due to their general nature, they lack of specific legalities required in the dynamic environment of the Generative AI and ML systems, which is essential today.
@@
Among the various neural network architectures, Convolutional Neural Networks (CNN) were the most familiar, with 7 out of 8 participants reporting having experience with them. Despite this familiarity, only 4 out of 8 participants perceived potential security issues in CNNs. In contrast, while only 1 participant had hands-on experience with Generative Adversarial Networks (GANs), a notable 4 participants identified security vulnerabilities in this model. This suggests that even limited experience with a specific neural network type can raise significant security concerns, possibly due to the inherent complexities or known vulnerabilities of that model. Expanding on this knowledge, the participants were also asked about the type of security issues they believe best describes the type of attack that could be done on neural networks. The majority (5 out of 8) of the participants believed that an attacker might be able to poison the training data by sending a bulk of malicious data to the model, while 2 out of 8 believed that an attacker might be able to modify input data and use that to force neural network make incorrect predictions, and also while only 1 out of 8 participants believed that an attacker might be able to reverse engineer the model itself and extract critical information.
@@ III-B3 Participant’s View: AML Experiences
Based on participant reports in this sample, we did not find evidence supporting H3. Moving forward, we asked participants about whether and how professionals in ML and security keep up to date with the latest security trends. 7 out of 12 stated that they actively keep themselves updated. Their primary reason to do so was professional growth in the industry and their secondary reasons were either due to company policies or personal interests in security. We also asked them about their sources to keep themselves updated. Some of them mentioned online resources such as courses related to AML or online blogs. Participant B8 mentioned “hacker news and feedly” and B3 mentioned “medium blogs”. However, social media platforms such as X (formerly Twitter) and LinkedIn are the most popular platforms for participants to stay updated about recent vulnerabilities and threats in the industry. Participant B12 stated that their knowledge was enhanced by mainstream media and related threat bulletins, while Participant B4 stated that they actually followed “online research papers from SaTML proceedings”. Notably, none of the participants mentioned CTF platforms as a source for staying updated on AML threats and defenses. In this sample, we did not find evidence supporting H2:
@@
Secondly, we asked participants about relevance of non-technical factors into adoption of defenses for AML, where we focused on three aspects—ethics, money, and convenience. This data highlights potential non-technical trade offs involved in adoption of AML techniques. Overall, ethics was the most valued, with a significant majority rating it as very or extremely important. Next was convenience (in the adoption of security practices in any ML model), which was rated important but less critical than ethics. Lastly, money, while being rated important by some participants, was not a necessary priority in this case. These responses are consistent with H4 in this sample:
@@ III-C Results
This study aimed to explore the different perspectives of 12 industry professionals specializing in security, ML and privacy. Our analysis confirmed that concerns about AML threats are significantly influenced by the participants’ background in security and education in privacy but not necessarily in ML. Contrary to our hypothesis, participants do not use Capture-the-Flag (CTF) platforms for any type of AML education, and instead preferred online resources and social media instead. The findings also revealed a lack of AML threat infrastructure and related defenses. Both technical factors, such as accuracy, time, and performance, and non-technical factors, including ethics, convenience, and financial considerations, were crucial in the adoption of AML security practices. Recommendations from participants included hiring security experts, providing technical training for ML engineers, implementing layered models, enhancing data processing, and adhering to better security hygiene practices to strengthen AML security.
@@ IV Study 2: A Capture-the-Flag Perspective
In this section, we explore an educational view of Adversarial Machine Learning threats. As seen in our previous studies, none of the participants mentioned CTFs as a learning resource for AML threats despite it being a popular source of cybersecurity education. In our second study, we explore whether CTF-style challenges can support learning about AML threats. For this study, our team developed CTF challenges for picoCTF wherein we developed Proof-of-Concept implementation of poisoning attack in the form of a chat bot. While the expression “gamification of security” can be a controversial one, CTF (an embodiment of this expression) is becoming more popular than ever. The study of Bratus et al. [6] explores a “Hacker Curriculum” and states that giving students a “security culture shock” using CTFs is important to help them question everything which is the ideal security professional mindset. Nelson et al. [30] designed a challenge-based curriculum where they concluded that despite the students finding the course pretty arduous, most of the students said they learned something new after solving those challenges. Similarly, in the studies of Dabrowski et al. [8] and Leune et al. [25], the competitive gaming-based approach was not only an effective educational tool amongst students but also increased their interest in cybersecurity. These studies provided the impetus for designing our novel Capture-the-Flag (CTF) challenges aimed at advancing education in adversarial machine learning (AML). Another paper [9], which aims to explore the role of red teaming in assessing Generative AI vulnerabilities, states that current red teaming exercises only serve as a regulatory check rather than a concrete security solution for Generative AI threats.
@@ IV-B1 Development of CTF Challenges
For this study, as mentioned above, we designed two Capture-the-Flag (CTF) challenges based on selected AML concepts. This experimental study included two sequenced tasks that were the CTF challenges developed by us and survey questions related to the participants’ background and the tasks. Each CTF challenge was time bounded to 15 minutes and the survey was time bounded to 30 minutes.
@@
Challenge 2 Overview: The second CTF challenge is a chatbot web application (Bake-time Lolly Chatbot), which is a website displaying different bakery products with an embedded chatbot that is used to engage users to help them with their orders. The development of this chatbot included the usage of the NTLK library [19] for text processing and a machine learning model using an SVM with sigmoid kernel [26]. In this case, the participant’s aim was to send malicious data into the chat bot in order to lower the confidence score and miss-classify the category of the product, which would poison the dataset to reveal a flag. (See the Appendix for more information.) The difficulty level of this challenge is assumed to be “advanced”. Participants are provided with the following interface and associated description:

### v2

@@ Abstract
An exponential growth of Machine Learning (ML) and its Generative AI applications brings with it significant security challenges, often referred to as Adversarial Machine Learning (AML). In this paper, we conducted two comprehensive studies to explore the perspectives of industry professionals and students on different AML vulnerabilities and their educational strategies. In our first study, we conducted an online survey with professionals revealing a notable correlation between cybersecurity education and concern for AML threats. For our second study, we developed two Capture-the-Flag (CTF) challenges that implement Natural Language Processing and Generative AI concepts and demonstrate a poisoning attack on the training dataset. The effectiveness of these challenges was evaluated by surveying undergraduate and graduate students at Carnegie Mellon University, finding that a CTF-based approach effectively engages interest in AML threats. Based on the responses of the participants in our research, we provide detailed recommendations emphasizing the critical need for integrated security education within the ML curriculum.
Index Terms:
adversarial machine learning, generative AI, machine learning security, cybersecurity education, capture-the-flag challenges, user study
@@ I Introduction
The modern-day computing ecosystem consists of a large volume and variety of data. This data can be viewed in different ways by different professionals. For cybersecurity professionals, data is primarily seen through the lens of security, encompassing access levels, authorization, and data integrity. In contrast, machine learning professionals approach data as a resource, primarily for analysis, focusing on datasets and models. To explore the similarities and differences in these views, as well as to address various aspects of security, privacy, and machine learning models in today’s world, we conducted comprehensive user studies. By capturing and analyzing these different perceptions, our goal was to bridge the gap between industry practices and academic research for AML threats, while developing a Capture-the-Flag (CTF) based educational approach for Adversarial Machine Learning (AML).
@@
In this paper, we undertake a thorough exploration of AML, examining various aspects which include theory on AML attacks and their implications, expanding our knowledge of AML attack to practical prompt-based learning models, a CTF perspective on AML, versus an industry perspective of AML. In this research, we explore different aspects of issues in Generative AI through the lens of two user studies and derive a novel educational approach for students and professionals for a defense strategy against AML attacks.
The first study involves an online survey of $12$ cybersecurity and machine learning experts where we collected and analyzed their experiences and perspectives on AML trends and defenses in the industry. In our second study, we used the concept of Generative AI to construct two CTF challenges which allowed us to evaluate two perspectives—develop our proof of concept for ML security CTF challenges for the picoCTF platform (a widely used university-hosted Capture-the-Flag educational website) and evaluate CTF as a mode of education for AML threats. In the same study, we also developed a CTF challenge where the user has to utilize Feature Collision and Convex Polytope attacks against a Generative AI chat bot with data loopback to influence the training data.
@@
  • We developed and evaluated our novel CTF challenges that are a proof-of-concept implementation of AML concepts, thus providing a practical and engaging learning approach for students.
@@ II-A Taxonomy of AML Attacks
Huang et al. [22] categorize AML attacks based on three properties:
@@
  • Oracle Attacks: In this type of attack, adversaries exploit the input-output pairings to gather information on the model like the parameters (Extraction attack), reconstruct training data (Inversion attacks) and determine in a brute force manner whether any of the supplied points were part of training dataset (Membership Inference Attack).
@@ II-B User-Centric Perspective of AML
Although there are various AML defense strategies and related research, we believe it is imperative to explore the impact and cause of human perception on the efficacy of these strategies. Not only does this point of view help devise effective defenses, it also helps us understand the best technique of learning for the people who are actually involved in this line of work.
@@
While the aforementioned studies give a deeper insight into AML and the need for its education, there seems to be little light on comparison of perspectives of two pivotal entities of the field: the professionals and students in the fields of Cybersecurity and Machine Learning. Hence, our aim is to identify the current attack landscape in the industry and associated challenges faced by students and professionals in detecting and mitigating AML attacks. We also aim to explore an academic perspective by developing AML CTF challenges in order to understand the students’ point of view for this new concept in cybersecurity.
@@ II-C Defenses against AML
In this section, we review prior research on defensive strategies against AML threats to understand the current security framework for ML systems. Goodfellow et al. [11] theorize a “Generative Adversarial Networks (GAN)” framework for estimating generative models using adversarial networks, and the authors proposed to train two models:
@@ III-A1 Ethical Considerations
Our data collection process for our study was organized in such a way that ethical considerations were taken into account in each step of the way. Our survey script, acknowledgments, recruitment materials, and other relevant documents were approved by our institution’s Institutional Review Board (IRB). The participants were compensated 5 USD for the 10-minute online survey, which is equivalent to 30 USD per hour, which is well above the minimum wage.
@@ III-A2 Recruitment
We conducted an online survey of 12 participants who are working professionals with varied backgrounds in Security, Machine Learning, and Privacy. To streamline participant selection, we utilized two platforms: WiCyS (Women in Cybersecurity Conference) 2023 [38] where participants were recruited through posters distributed to attendees, and an online platform connecting students of our university with their alumni where we post our recruitment text on a message board. The interested participants in both cases were asked to contact us by email. After receiving 24 inquiries collectively from both platforms, we evaluated participants’ backgrounds to verify that they were above the age of 18 and that they are a working professionals in either ML or security, and selected 12 qualifying participants out of them. The survey was hosted on Qualtrics [31] and each participant was sent a unique link to record their responses.
@@ III-B Analysis
Our survey was taken by 12 participants. About 64% of the participants belonged to an age ranging between 25 and 34 years. The data collected was checked for possible errors, such as missing or irrelevant text entries. We also checked responses to see if any participants had zero experience in all three fields—Cybersecurity, Privacy, and Machine Learning. After finding no errors, we analyzed the study data as listed below.
@@ III-B1 Background of Participants
Our understanding of the correlation between the participants’ background in Machine Learning, Cybersecurity and Privacy and their concern for security in Artificial Intelligence and Machine Learning led us to ask them non-personally identifiable information such as participants’ age, gender, education level and relevant experience in the above-mentioned fields. Regarding the background of the participants, we analyzed two key metrics, namely “Education” and “Work Experience”. The table below shows the level of education and the level of experience that correspond to the number of participants in the given category. This table (Table I) shows a categorization as relevant to our hypothesis by aggregating some results. The more detailed breakdown is available in the Appendix.
@@
As seen in the table, the participants in this study display a diverse background and a set of qualified skills, demonstrating proficiency in matters related to security and privacy issues in Machine Learning models. About 95% of the participants have worked with the ML algorithms in some form—Regression models (9 out of 12), Decision trees (10 out of 12), Support Vector Machines (6 out of 12), Neural Networks (9 out of 12) and others including Clustering algorithms (1 out of 12) and reinforcement learning (1 out of 12). Based on these observations, we explore our first hypothesis:
@@
From the above table, we observe an “expertise effect” for participants having a background in education or work experience in cybersecurity since their backgrounds are correlated with their concern for AML threats, and hence are more likely to think about security in ML. However, a background in ML itself does not show any correlation with the participant’s concern for security issues in ML. Lastly, an educational background in privacy is correlated with the participant’s concern, whereas merely work experience is not. This result highlights the importance of security and privacy education amongst professionals, especially ML engineers.
@@ III-B2 Personal Experience in the industry
In the studies of Hickman et al. [20] and Rees et al. [32], the authors reviewed official documents, which consisted of EU guidelines for Trustworthy AI and organizational and government documents, respectively. In both cases, the authors have concluded that while the documents aim to address the concerns in AI, due to their general nature, they lack specific legalities required in the dynamic environment of the Generative AI and ML systems, which is essential today.
@@
Among the various neural network architectures, Convolutional Neural Networks (CNN) were the most familiar, with 7 out of 8 participants reporting having experience with them. Despite this familiarity, only 4 out of 8 participants perceived potential security issues in CNNs. In contrast, while only 1 participant had hands-on experience with Generative Adversarial Networks (GANs), a notable 4 participants identified security vulnerabilities in this model. This suggests that even limited experience with a specific neural network type can raise significant security concerns, possibly due to the inherent complexities or known vulnerabilities of that model. Expanding on this knowledge, the participants were also asked about the type of security issues they believe best describes the type of attack that could be done on neural networks. The majority (5 out of 8) of the participants believed that an attacker might be able to poison the training data by sending a bulk of malicious data to the model, while 2 out of 8 believed that an attacker might be able to modify input data and use that to force a neural network to make incorrect predictions, and also while only 1 out of 8 participants believed that an attacker might be able to reverse engineer the model itself and extract critical information.
@@ III-B3 Participant’s View: AML Experiences
Based on participant reports in this sample, we did not find evidence supporting H3. Moving forward, we asked participants about whether and how professionals in ML and security keep up to date with the latest security trends. 7 out of 12 stated that they actively keep themselves updated. Their primary reason to do so was professional growth in the industry and their secondary reasons were either due to company policies or personal interests in security. We also asked them about their sources to keep themselves updated. Some of them mentioned online resources such as courses related to AML or online blogs. Participant B8 mentioned “hacker news and feedly” and B3 mentioned “medium blogs”. However, social media platforms such as X (formerly Twitter) and LinkedIn are the most popular platforms for participants to stay updated about recent vulnerabilities and threats in the industry. Participant B12 stated that their knowledge was enhanced by mainstream media and related threat bulletins, while Participant B4 stated that they actually followed “online research papers from Security and Trustworthy Machine Learning (SaTML) proceedings”. Notably, none of the participants mentioned CTF platforms as a source for staying updated on AML threats and defenses. In this sample, we did not find evidence supporting H2:
@@
Secondly, we asked participants about relevance of non-technical factors into adoption of defenses for AML, where we focused on three aspects—ethics, money, and convenience. This data highlights potential non-technical trade-offs involved in adoption of AML techniques. Overall, ethics was the most valued, with a significant majority rating it as very or extremely important. Next was convenience (in the adoption of security practices in any ML model), which was rated important but less critical than ethics. Lastly, money, while being rated important by some participants, was not a necessary priority in this case. These responses are consistent with H4 in this sample:
@@ III-C Results
This study aimed to explore the different perspectives of 12 industry professionals specializing in security, ML and privacy. Our analysis confirmed that concerns about AML threats are significantly influenced by the participants’ background in security and education in privacy but not necessarily in ML. Contrary to our hypothesis, participants do not use CTF platforms for any type of AML education, and instead preferred online resources and social media. The findings also revealed a lack of AML threat infrastructure and related defenses. Both technical factors, such as accuracy, time, and performance, and non-technical factors, including ethics, convenience, and financial considerations, were crucial in the adoption of AML security practices. Recommendations from participants included hiring security experts, providing technical training for ML engineers, implementing layered models, enhancing data processing, and adhering to better security hygiene practices to strengthen AML security.
@@ IV Study 2: A Capture-the-Flag Perspective
In this section, we explore an educational view of Adversarial Machine Learning threats. As seen in our previous studies, none of the participants mentioned CTFs as a learning resource for AML threats despite it being a popular source of cybersecurity education. In our second study, we explore whether CTF-style challenges can support learning about AML threats. For this study, our team developed CTF challenges for picoCTF wherein we developed a Proof-of-Concept implementation of a poisoning attack in the form of a chat bot. While the expression “gamification of security” can be a controversial one, CTF (an embodiment of this expression) is becoming more popular than ever. The study of Bratus et al. [6] explores a “Hacker Curriculum” and states that giving students a “security culture shock” using CTFs is important to help them question everything which is the ideal security professional mindset. Nelson et al. [30] designed a challenge-based curriculum where they concluded that despite the students finding the course pretty arduous, most of the students said they learned something new after solving those challenges. Similarly, in the studies of Dabrowski et al. [8] and Leune et al. [25], the competitive gaming-based approach was not only an effective educational tool amongst students but also increased their interest in cybersecurity. These studies provided the impetus for designing our novel CTF challenges aimed at advancing education in AML. Another paper [9], which aims to explore the role of red teaming in assessing Generative AI vulnerabilities, states that current red teaming exercises only serve as a regulatory check rather than a concrete security solution for Generative AI threats.
@@ IV-B1 Development of CTF Challenges
For this study, as mentioned above, we designed two CTF challenges based on selected AML concepts. This experimental study included two sequenced tasks that were the CTF challenges developed by us and survey questions related to the participants’ background and the tasks. Each CTF challenge was time bounded to 15 minutes and the survey was time bounded to 30 minutes.
@@
Challenge 2 Overview: The second CTF challenge is a chatbot web application (Bake-time Lolly Chatbot), which is a website displaying different bakery products with an embedded chatbot that is used to engage users to help them with their orders. The development of this chatbot included the usage of the NLTK library [19] for text processing and a machine learning model using an Support Vector Machine (SVM) with sigmoid kernel [26]. In this case, the participant’s aim was to send malicious data into the chat bot in order to lower the confidence score and misclassify the category of the product, which would poison the dataset to reveal a flag. (See the Appendix for more information.) The difficulty level of this challenge is assumed to be “advanced”. Participants are provided with the following interface and associated description: